### v1

@@ 4.3 Variance Estimation
We establish the consistency of this estimator. Because the estimators contain averages over random functions, e.g. $\widehat{k}^{(\tau)}_{1}(x,y)$, and the asymptotic (co)variances featuring $\rho$ even contain averages of averages of random functions, the consistency is not trivial. Yet, it follows by standard arguments on random functions from empirical process theory using the Glivenko-Cantelli theorem. For details, see the proof in Appendix B and vandervaart2000.
@@ 4.4 Confidence Intervals, Tests, and Fisher Transformation
It is well known that for the empirical Pearson correlation, the normal limiting distribution does not yield a good approximation if the true Pearson correlation is close to 1-1 or $1$, because then the true sampling distribution is heavily skewed. A solution to this problem is to transform the empirical correlation to an unbounded scale (the real line) by the Fisher transformation (fisher1915), and to use the limiting distribution of the Fisher transformation instead. This yields a much better approximation close to the boundaries and leaves the situation essentially unchanged away from the boundaries. This reasoning carries over to other correlation coefficients as well, see pohle2024measuring for a detailed discussion of the Fisher transformation for generic dependence measures. The Fisher transformation is defined as
@@ 4.5 Asymptotic Distributions under Independence
Corollary 4.5 nicely extends the classical results for the continuous case, where the asymptotic variances are $4/9$ and $1$, respectively, to the general case, where the additional factors involving the double-tie probabilities of $X$ and $Y$ arise. Even though the asymptotic variances are quite fundamental and well known for the continuous case, see kendall1938 and hoeffding1948, we are not aware of any reference for their neat formulas in the general case. Interestingly, already hoeffding1948 conjectured that they depend on the tie probabilities $\mathrm{P}(X=X^{\prime})$ and $\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$.
Note that by the equality of $k_{1}^{(\tau)}$ and $k_{1}^{(\rho)}$, and our multivariate CLT for U-statistics (Proposition A.3 in the appendix), the asymptotic correlation between $\widehat{\tau}$ and $\widehat{\rho}$ under independence is 1. Considering the ratio of the standard deviations, this implies that they asymptotically have the deterministic relationship $\widehat{\rho}=3/2\,\widehat{\tau}$, which was already proven by daniels1944. Thus, the higher asymptotic variance of $\widehat{\rho}$ is not caused by it being less efficient, but by it taking higher (absolute) values under independence.
@@ 5.1 Consistency
Absolute regularity or $\beta$-mixing is a form of asymptotic independence. It is stronger than strong mixing and weaker than uniform mixing, see bradley2005 for details.
@@
Strong consistency of all five empirical rank correlations follows directly by an appropriate law of large numbers for U-statistics under weak dependence (dehling2006, Theorem 1), together with the continuous mapping theorem in the cases of $\widehat{\gamma}$, $\widehat{\tau}_{b}$, and $\widehat{\rho}_{b}$.
@@ 5.2 Asymptotic Distributions
To generalize the results from Propositions 4.2 and 4.3, we need a univariate central limit theorem under Assumption 2 (see denker1983 and Proposition A.2 in the appendix), a multivariate version of it (Proposition A.3 in the appendix), and again the delta method. Furthermore, for the derivation of the asymptotic variances, we need Lemma B.2 in the appendix.
@@
Thus, only the expressions for the asymptotic variances and covariances of the U-statistics occurring in the definitions of the rank correlations change compared to the iid case. In the continuous case, dehling2017 obtained the limiting distribution for Kendall’s $\tau$ under similar assumptions.
@@
Such an absolute summability condition on the autocovariances is very common as an assumption in time series analysis, ruling out long memory (e.g. hassler2018). Here, it is already implied by Assumption 2 and the boundedness of our kernels.
@@ 5.3 Variance Estimation
Compared to the estimator $\widehat{\sigma}_{lm,\textup{iid}}$ from Subsection 4.3, the additional difficulty arises from the infinite sum in (22). Besides estimating the functions $k_{1}^{(l)}$ and $k_{1}^{(m)}$ as laid out in Subsection 4.3, we now employ classical techniques for long-run variance estimation, which are well known from the case of the sample mean (newey1987). That is, we estimate $\sigma_{lm}$ via
@@
We refer to lazarus2018 for an overview of the vast literature on long-run variance estimation, or heteroscedastiticy and autocorrelation consistent (HAC) estimation, in the case of sample means or regression coefficients. dehling2017 propose the use of a HAC estimator of this type for U-statistics of order 2 and especially for Kendall’s $\tau$, but under a different set of assumptions. We establish consistency of our variance estimators under our assumptions on the time series and standard assumptions on the kernel and the bandwidth.
@@
We use the Bartlett kernel, or triangular kernel, which has been popularized by newey1987:
@@
As the bandwidth, we choose $b_{n}=\lfloor 2n^{1/3}\rfloor$ as in dehling2017. With these choices, Assumption 3 is satisfied. We conjecture that the bounded support assumption for the kernel can be replaced by an assumption on its rate of decay.
It is well known that variance estimation under temporal dependence is a notoriously difficult problem, usually plagued by oversized tests and confidence intervals with undercoverage, with the problem getting worse with the strength of temporal dependence and only disappearing for large sample sizes, see again lazarus2018 on this issue. Alternatives to the HAC estimator such as the moving block bootstrap also exhibit this behaviour (fitzenberger1998). As expected, these problems also arise in our simulations, see Section 6.
@@ 5.4 Variance Formulas under Independence
Note that in order to achieve compact formulas, we have expressed the asymptotic variances for $\tau,\rho,\gamma$ and $\tau_{b}$ via the autocorrelation in terms of $\rho$, and the asymptotic variance for the grade correlation via the autocorrelation in terms of $\rho_{b}$. We could also express all results in terms of one or the other, see Definition 2.6 and Lemma B.4 in the appendix for their relation. In the continuous case, this result for $\widehat{\tau}$ and $\widehat{\rho}$ has recently been derived by lun2023.
@@ 7.1 Ecological Data
We consider the bivariate count data $(x_{1},y_{1}),\ldots,(x_{100},y_{100})$ in Table 1 of holgate66, which were further analyzed by kokonendji18. The data arose in a study of a secondary rain forest in Trinidad, and they express the numbers of plants of the species Lacistema aggregatum ($x_{i}$) and Protium guianense ($y_{i}$) in each of 100 contiguous quadrants. Figure 2 visualizes the data using a bubble plot.
@@ 7.2 Accidents Data
Our second example is about bivariate time series data, namely the bivariate counts $(x_{i},y_{i})$, $i=1,\ldots,365$, which refer to the daily numbers of daytime ($x_{i}$) and nighttime ($y_{i}$) road accidents in the Schiphol area (Netherlands) during the year 2001. The data have been read from Figure 1 in pedeli11. These authors showed that the data exhibit significant serial dependence and propose to model the data by a bivariate INAR(1)(1) model. In Figure 3 we again depict the empirical joint distribution in a bubble plot. Figure 4 presents a time plot of both series.
@@
Here, our interest is in the cross-dependence between the daytime and nighttime accidents, where hypothesis tests and confidence intervals have to account for the serial dependence (without imposing a specific time series model) as described in Section 5. All rank correlations and Pearson correlation indicate mild positive positive dependence (see Table 2) with values around 0.1. Interestingly and contrary to the first data example, Pearson correlation is slightly larger than all of the rank correlations here. Usually, Pearson correlation being larger than Kendall’s $\tau$ or Spearman’s $\rho$ is associated with linear dependence, that is, a form of dependence that can be captured well by Pearson correlation. A more in-depth explanation is as follows: The shape of the bubble plot suggests that the joint distribution can be well approximated by an elliptical distribution. For elliptical distributions, firstly, the dependence is fully characterized by Pearson correlation $r$, and secondly, there are well-known functional relationships between $r$ and the classical rank correlations $\tau$ and $\rho$, which show that these rank correlations are always smaller than $r$ (embrechts2002). The respective generalized rank correlations are greater than $\tau$ and $\rho$, but still smaller than $r$ here.
@@ 8 Conclusion
Although asymptotic confidence intervals and tests are the most classical approach to statistical inference, and Kendall’s $\tau$ and Spearman’s $\rho$ have favorable properties compared to Pearson correlation and are commonly employed in descriptive statistics, it is quite surprising that asymptotic inference for these rank correlations has been lacking fundamental results and, thus, has rarely been used in inductive statistics. Consequently, sound inferential procedures (with the exception of independence tests) are not implemented in most statistical software. The results of this paper, which of course build on earlier foundational work such as hoeffding1948 and dehling2017, allow for a change in this respect.
@@
For time series data, we discover some problems with variance estimation under strong temporal dependence and in small to medium-sized samples. This, however, is not surprising as variance estimation under temporal dependence in similar settings (e.g. for the sample mean or linear regression coefficients) is a notoriously difficult problem, plagued by oversized tests and confidence intervals with undercoverage. Despite these issues, we consider our asymptotic confidence intervals and tests to be a substantial improvement on the status quo in time series settings, where inferential procedures are hardly available so far. Future research in this direction is undoubtedly crucial. For example, self-normalization might be a way to tackle these problems (shao2015), or at least the improvement of the HAC estimator by systematically comparing different choices of the bandwidth or kernel.
While Kendall’s $\tau$, Spearman’s $\rho$, and their variants for the discrete case, are the classical and most widely-used rank correlations, inference for other representatives of this class is also relevant and can be approached by using the results of this paper. For example, there are asymmetric versions of $\tau_{b}$ and grade correlation: Somers’ $D$ defined as $\tau(X,Y)/\tau(X,X)$ (somers1962; newson2002), and the recently introduced asymmetric grade correlation defined as $\rho(X,Y)/\rho(X,X)$ (walz2025), measure directed dependence of $Y$ on $X$, e.g. to assess the predictive potential. In the continuous case, they equal $\tau$ and $\rho$, respectively, and for a dichotomous $Y$, they are both related by a one-to-one mapping to the area under the receiver operating characteristic (ROC) curve, short area under the curve (AUC), which is a popular measure to assess the performance of binary classifiers (walz2025). For all these measures, inference tools can be constructed by using the techniques introduced in this paper. walz2025 do this for the case of asymmetric grade correlation and develop tests for the equality of two grade correlations as well. Spearman’s $\rho$ and asymmetric grade correlation are also closely related to the slope coefficients of rank-rank regressions, for which inference has recently been treated by chetverikov2023.
[…]
@@ Appendix A Asymptotic Theory for U-Statistics under Weak Dependence
A U-statistic can be viewed as a generalized sample mean, namely as an average over a function of $r$ sample elements instead of only one element. See e.g. lee1990 for an introduction to U-statistics, and dehling2006 for a discussion of limit theorems for U-statistics under weak dependence. Let $\{\mathbf{X}_{i}\}_{i\in\mathbb{Z}}$ be a $K$-variate stochastic process, $\mathbf{X}_{i}^{\prime}=\left(X_{1,t},\ldots,X_{K,t}\right)$, and $\{\mathbf{X}_{i}\}_{i=1}^{n}$ a sample of size $n$ arising from it, or in other words, a time series of length $n$. By $\mathbf{X}^{(2)}_{i},\mathbf{X}^{(3)}_{i}$$\ldots,$ we denote independent copies of $\mathbf{X}_{i}$, where $\mathbf{X}_{i}$ itself is also written as $\mathbf{X}^{(1)}_{i}$ in this context.
@@
$U_{\mathbf{X}}(k)=\binom{n}{r}^{-1}\sum_{1\leq t_{1}<\ldots<t_{r}\leq n}k\left(\mathbf{X}_{t_{1}},\ldots,\mathbf{X}_{t_{r}}\right),$
@@ A U-statistic is an unbiased estimator for the parameter
We usually leave out the subscript $\mathbf{X}$} in $U_{\mathbf{X}}$ if it is not necessary to emphasize based on which process the U-statistic is defined, and analogously the argument $k$ if the same holds true for the kernel. Furthermore, we omit the time index $t$ in situations where we do not need to distinguish different points in time.
@@
A key tool for the analysis of U-statistics is the Hoeffding decomposition (hoeffding1961). For our purposes, we focus on the leading term, as the lower-order terms do not play a role in the asymptotic analysis, and we consider an asymptotic version of the decomposition: A U-statistic can be rewritten as (under our assumptions spelled out in Proposition A.2 below)
@@
This decomposition shows that the U-statistic $U$ can essentially be treated as a sample mean of $k_{1}(\mathbf{X}_{i})$. This is the essence behind central limit theorems for U-statistics, which consequently can be derived from usual central limit theorems. The kernel $k_{1}$ (derived from $k$ and the distribution of $(\mathbf{X}^{(2)},\ldots\mathbf{X}^{(r)})$) is of order 1 and (amongst others) plays a central role in determining limiting variances in central limit theorems for U-statistics. The following central limit theorem is due to denker1983. We restrict attention to the case of bounded kernels, which we exclusively deal with in this paper. Boundedness could be replaced by a moment condition at the cost of strengthening the mixing condition.
@@
The asymptotic Hoeffding decomposition (A.3) explains the formula for the asymptotic variance $\sigma_{U}^{2}$. It has the form of the asymptotic variance of the process $\{k_{1}(\mathbf{X}_{i})\}_{i\in\mathbb{Z}}$, i.e. the variance of the sample mean (scaled by $\sqrt{n}$) of this process, which is a long-run variance matrix in the time series case. Note that if the kernel $k_{1}$ is degenerate, i.e. $\mathrm{Var}(k_{1}(X_{i}))=0$, the limiting distribution becomes degenerate, so a stronger scaling factor than $\sqrt{n}$ is needed to obtain a non-degenerate limiting distribution. But this is not relevant for our paper as the kernels showing up here are only degenerate in pathological cases, see hoeffding1948; dengler2010 for a discussion in the case of $\widehat{\tau}$.
There is only a small literature on variance estimation for U-statistics: For the iid case, see shirahata1992, wang2014 and iwashita2024; for the time series case the only paper that we are aware of is dehling2017, which focuses on U-statistics of order two.
We finally provide a multivariate central limit theorem, which is the basis for many results in this paper and which follows from the univariate version in Proposition A.2 by the Cramér-Wold device and a combinatorial argument. We consider a vector of U-statistics and allow them to depend on different underlying processes. For a multivariate central limit theorem for U-statistics in the iid case see hoeffding1948.
@@
Let $\{(X_{1,t},\ldots,X_{K,t})\}_{i\in\mathbb{Z}}$ be a $K$-dimensional process, where all components satisfy Assumption 1 or Assumption 2. Let $\mathbf{U}=(U^{(1)}_{\mathbf{X}_{I^{(1)}}}(k^{(1)}),\ldots,U^{(D)}_{\mathbf{X}_{I^{(D)}}}(k^{(D)}))^{\prime}$ be a $D$-dimensional vector of U-statistics with bounded symmetric kernels $\mathbf{k}=(k^{(1)},\ldots,k^{(D)})^{\prime}$ of orders $(r^{(1)},\ldots,r^{(D)})^{\prime}$ and underlying stochastic processes $\{(\mathbf{X}_{I^{(1)},t},\ldots,\mathbf{X}_{I^{(D)},t})\}_{i\in\mathbb{Z}}$, where $I^{(d)}\subset\{1,\ldots,K\}$ for $d=1,\ldots,D$ are subsets of the original index set. Furthermore, let $\theta^{(d)}$ and $k_{1}^{(d)}$, $d=1,\ldots,D$, denote the quantities defined in (A.1) and (A.4), respectively, for the $d$-th U-statistic, and define $\mbox{$\theta$}:=(\theta^{(1)},\ldots,\theta^{(D)})^{\prime}$. Then it holds that
@@
$\sigma_{vw}=\mathrm{E}\big[k^{(v)}_{1}(\mathbf{X}_{I^{(v)},t})\,k^{(w)}_{1}(\mathbf{X}_{I^{(w)},t})\big],$
@@
$\sigma_{vw}=\sum_{h=-\infty}^{\infty}\mathrm{E}\big[k^{(v)}_{1}(\mathbf{X}_{I^{(v)},t})\,k^{(w)}_{1}(\mathbf{X}_{I^{(w)},t+h})\big].$
@@
$U^{(d)}_{\mathbf{X}_{I^{(d)}}}(k^{(d)})=\binom{n}{r^{(d)}}^{-1}\sum_{1\leq t_{1}<\ldots<t_{r^{(d)}}\leq n}k^{(d)}\left(\mathbf{X}_{t_{1}},\ldots,\mathbf{X}_{t_{r^{(d)}}}\right)$

### v2

@@ 4.3 Variance Estimation
We establish the consistency of this estimator. Because the estimators contain averages over random functions, e.g. $\widehat{k}^{(\tau)}_{1}(x,y)$, and the asymptotic (co)variances featuring $\rho$ even contain averages of averages of random functions, the consistency is not trivial. Yet, it follows by standard arguments on random functions from empirical process theory using the Glivenko-Cantelli theorem. For details, see the proof in Appendix B and Van der Vaart (2000, Chapter 19.4).
@@ 4.4 Confidence Intervals, Tests, and Fisher Transformation
It is well known that for the empirical Pearson correlation, the normal limiting distribution does not yield a good approximation if the true Pearson correlation is close to 1-1 or $1$, because then the true sampling distribution is heavily skewed. A solution to this problem is to transform the empirical correlation to an unbounded scale (the real line) by the Fisher transformation (Fisher, 1915), and to use the limiting distribution of the Fisher transformation instead. This yields a much better approximation close to the boundaries and leaves the situation essentially unchanged away from the boundaries. This reasoning carries over to other correlation coefficients as well, see Pohle et al. (2024, Appendix B) for a detailed discussion of the Fisher transformation for generic dependence measures. The Fisher transformation is defined as
@@ 4.5 Asymptotic Distributions under Independence
Corollary 4.5 nicely extends the classical results for the continuous case, where the asymptotic variances are $4/9$ and $1$, respectively, to the general case, where the additional factors involving the double-tie probabilities of $X$ and $Y$ arise. Even though the asymptotic variances are quite fundamental and well known for the continuous case, see Kendall (1938) and Hoeffding (1948), we are not aware of any reference for their neat formulas in the general case. Interestingly, already Hoeffding (1948, p. 317, 320) conjectured that they depend on the tie probabilities $\mathrm{P}(X=X^{\prime})$ and $\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$.
Note that by the equality of $k_{1}^{(\tau)}$ and $k_{1}^{(\rho)}$, and our multivariate CLT for U-statistics (Proposition A.3 in the appendix), the asymptotic correlation between $\widehat{\tau}$ and $\widehat{\rho}$ under independence is 1. Considering the ratio of the standard deviations, this implies that they asymptotically have the deterministic relationship $\widehat{\rho}=3/2\,\widehat{\tau}$, which was already proven by Daniels (1944). Thus, the higher asymptotic variance of $\widehat{\rho}$ is not caused by it being less efficient, but by it taking higher (absolute) values under independence.
@@ 5.1 Consistency
Absolute regularity or $\beta$-mixing is a form of asymptotic independence. It is stronger than strong mixing and weaker than uniform mixing, see Bradley (2005) for details.
@@
Strong consistency of all five empirical rank correlations follows directly by an appropriate law of large numbers for U-statistics under weak dependence (Dehling, 2006, Theorem 1), together with the continuous mapping theorem in the cases of $\widehat{\gamma}$, $\widehat{\tau}_{b}$, and $\widehat{\rho}_{b}$.
@@ 5.2 Asymptotic Distributions
To generalize the results from Propositions 4.2 and 4.3, we need a univariate central limit theorem under Assumption 2 (see Denker and Keller (1983) and Proposition A.2 in the appendix), a multivariate version of it (Proposition A.3 in the appendix), and again the delta method. Furthermore, for the derivation of the asymptotic variances, we need Lemma B.2 in the appendix.
@@
Thus, only the expressions for the asymptotic variances and covariances of the U-statistics occurring in the definitions of the rank correlations change compared to the iid case. In the continuous case, Dehling et al. (2017) obtained the limiting distribution for Kendall’s $\tau$ under similar assumptions.
@@
Such an absolute summability condition on the autocovariances is very common as an assumption in time series analysis, ruling out long memory (e.g. Hassler, 2018). Here, it is already implied by Assumption 2 and the boundedness of our kernels.
@@ 5.3 Variance Estimation
Compared to the estimator $\widehat{\sigma}_{lm,\textup{iid}}$ from Subsection 4.3, the additional difficulty arises from the infinite sum in (22). Besides estimating the functions $k_{1}^{(l)}$ and $k_{1}^{(m)}$ as laid out in Subsection 4.3, we now employ classical techniques for long-run variance estimation, which are well known from the case of the sample mean (Newey and West, 1987). That is, we estimate $\sigma_{lm}$ via
@@
We refer to Lazarus et al. (2018) for an overview of the vast literature on long-run variance estimation, or heteroscedastiticy and autocorrelation consistent (HAC) estimation, in the case of sample means or regression coefficients. Dehling et al. (2017) propose the use of a HAC estimator of this type for U-statistics of order 2 and especially for Kendall’s $\tau$, but under a different set of assumptions. We establish consistency of our variance estimators under our assumptions on the time series and standard assumptions on the kernel and the bandwidth.
@@
We use the Bartlett kernel, or triangular kernel, which has been popularized by Newey and West (1987):
@@
As the bandwidth, we choose $b_{n}=\lfloor 2n^{1/3}\rfloor$ as in Dehling et al. (2017). With these choices, Assumption 3 is satisfied. We conjecture that the bounded support assumption for the kernel can be replaced by an assumption on its rate of decay.
It is well known that variance estimation under temporal dependence is a notoriously difficult problem, usually plagued by oversized tests and confidence intervals with undercoverage, with the problem getting worse with the strength of temporal dependence and only disappearing for large sample sizes, see again Lazarus et al. (2018) on this issue. Alternatives to the HAC estimator such as the moving block bootstrap also exhibit this behaviour (Fitzenberger, 1998). As expected, these problems also arise in our simulations, see Section 6.
@@ 5.4 Variance Formulas under Independence
Note that in order to achieve compact formulas, we have expressed the asymptotic variances for $\tau,\rho,\gamma$ and $\tau_{b}$ via the autocorrelation in terms of $\rho$, and the asymptotic variance for the grade correlation via the autocorrelation in terms of $\rho_{b}$. We could also express all results in terms of one or the other, see Definition 2.6 and Lemma B.4 in the appendix for their relation. In the continuous case, this result for $\widehat{\tau}$ and $\widehat{\rho}$ has recently been derived by Lun et al. (2023, Corollary 2.1).
@@ 7.1 Ecological Data
We consider the bivariate count data $(x_{1},y_{1}),\ldots,(x_{100},y_{100})$ in Table 1 of Holgate (1966), which were further analyzed by Kokonendji and Puig (2018). The data arose in a study of a secondary rain forest in Trinidad, and they express the numbers of plants of the species Lacistema aggregatum ($x_{i}$) and Protium guianense ($y_{i}$) in each of 100 contiguous quadrants. Figure 2 visualizes the data using a bubble plot.
@@ 7.2 Accidents Data
Our second example is about bivariate time series data, namely the bivariate counts $(x_{i},y_{i})$, $i=1,\ldots,365$, which refer to the daily numbers of daytime ($x_{i}$) and nighttime ($y_{i}$) road accidents in the Schiphol area (Netherlands) during the year 2001. The data have been read from Figure 1 in Pedeli and Karlis (2011). These authors showed that the data exhibit significant serial dependence and propose to model the data by a bivariate INAR(1)(1) model. In Figure 3 we again depict the empirical joint distribution in a bubble plot. Figure 4 presents a time plot of both series.
@@
Here, our interest is in the cross-dependence between the daytime and nighttime accidents, where hypothesis tests and confidence intervals have to account for the serial dependence (without imposing a specific time series model) as described in Section 5. All rank correlations and Pearson correlation indicate mild positive positive dependence (see Table 2) with values around 0.1. Interestingly and contrary to the first data example, Pearson correlation is slightly larger than all of the rank correlations here. Usually, Pearson correlation being larger than Kendall’s $\tau$ or Spearman’s $\rho$ is associated with linear dependence, that is, a form of dependence that can be captured well by Pearson correlation. A more in-depth explanation is as follows: The shape of the bubble plot suggests that the joint distribution can be well approximated by an elliptical distribution. For elliptical distributions, firstly, the dependence is fully characterized by Pearson correlation $r$, and secondly, there are well-known functional relationships between $r$ and the classical rank correlations $\tau$ and $\rho$, which show that these rank correlations are always smaller than $r$ (Embrechts et al., 2002). The respective generalized rank correlations are greater than $\tau$ and $\rho$, but still smaller than $r$ here.
@@ 8 Conclusion
Although asymptotic confidence intervals and tests are the most classical approach to statistical inference, and Kendall’s $\tau$ and Spearman’s $\rho$ have favorable properties compared to Pearson correlation and are commonly employed in descriptive statistics, it is quite surprising that asymptotic inference for these rank correlations has been lacking fundamental results and, thus, has rarely been used in inductive statistics. Consequently, sound inferential procedures (with the exception of independence tests) are not implemented in most statistical software. The results of this paper, which of course build on earlier foundational work such as Hoeffding (1948) and Dehling et al. (2017), allow for a change in this respect.
@@
For time series data, we discover some problems with variance estimation under strong temporal dependence and in small to medium-sized samples. This, however, is not surprising as variance estimation under temporal dependence in similar settings (e.g. for the sample mean or linear regression coefficients) is a notoriously difficult problem, plagued by oversized tests and confidence intervals with undercoverage. Despite these issues, we consider our asymptotic confidence intervals and tests to be a substantial improvement on the status quo in time series settings, where inferential procedures are hardly available so far. Future research in this direction is undoubtedly crucial. For example, self-normalization might be a way to tackle these problems (Shao, 2015), or at least the improvement of the HAC estimator by systematically comparing different choices of the bandwidth or kernel.
While Kendall’s $\tau$, Spearman’s $\rho$, and their variants for the discrete case, are the classical and most widely-used rank correlations, inference for other representatives of this class is also relevant and can be approached by using the results of this paper. For example, there are asymmetric versions of $\tau_{b}$ and grade correlation: Somers’ $D$ defined as $\tau(X,Y)/\tau(X,X)$ (Somers, 1962; Newson, 2002), and the recently introduced asymmetric grade correlation defined as $\rho(X,Y)/\rho(X,X)$ (Walz et al., 2025), measure directed dependence of $Y$ on $X$, e.g. to assess the predictive potential. In the continuous case, they equal $\tau$ and $\rho$, respectively, and for a dichotomous $Y$, they are both related by a one-to-one mapping to the area under the receiver operating characteristic (ROC) curve, short area under the curve (AUC), which is a popular measure to assess the performance of binary classifiers (Walz et al., 2025). For all these measures, inference tools can be constructed by using the techniques introduced in this paper. Walz et al. (2025) do this for the case of asymmetric grade correlation and develop tests for the equality of two grade correlations as well. Spearman’s $\rho$ and asymmetric grade correlation are also closely related to the slope coefficients of rank-rank regressions, for which inference has recently been treated by Chetverikov and Wilhelm (2025).
@@ Appendix A Asymptotic Theory for U-Statistics under Weak Dependence
A U-statistic can be viewed as a generalized sample mean, namely as an average over a function of $r$ sample elements instead of only one element. See e.g. Lee (1990) for an introduction to U-statistics, and Dehling (2006) for a discussion of limit theorems for U-statistics under weak dependence. Let $\{\mathbf{X}_{i}\}_{i\in\mathbb{Z}}$ be a $K$-variate stochastic process, $\mathbf{X}_{i}^{\prime}=\left(X_{1,i},\ldots,X_{K,i}\right)$, and $\{\mathbf{X}_{i}\}_{i=1}^{n}$ a sample of size $n$ arising from it, or in other words, a time series of length $n$. By $\mathbf{X}^{(2)}_{i},\mathbf{X}^{(3)}_{i}$$\ldots,$ we denote independent copies of $\mathbf{X}_{i}$, where $\mathbf{X}_{i}$ itself is also written as $\mathbf{X}^{(1)}_{i}$ in this context.
@@
$U_{\mathbf{X}}(k)=\binom{n}{r}^{-1}\sum_{1\leq i_{1}<\ldots<i_{r}\leq n}k\left(\mathbf{X}_{i_{1}},\ldots,\mathbf{X}_{i_{r}}\right),$
@@ A U-statistic is an unbiased estimator for the parameter
We usually leave out the subscript $\mathbf{X}$} in $U_{\mathbf{X}}$ if it is not necessary to emphasize based on which process the U-statistic is defined, and analogously the argument $k$ if the same holds true for the kernel. Furthermore, we omit the time index $i$ in situations where we do not need to distinguish different points in time.
@@
A key tool for the analysis of U-statistics is the Hoeffding decomposition (Hoeffding, 1961). For our purposes, we focus on the leading term, as the lower-order terms do not play a role in the asymptotic analysis, and we consider an asymptotic version of the decomposition: A U-statistic can be rewritten as (under our assumptions spelled out in Proposition A.2 below)
@@
This decomposition shows that the U-statistic $U$ can essentially be treated as a sample mean of $k_{1}(\mathbf{X}_{i})$. This is the essence behind central limit theorems for U-statistics, which consequently can be derived from usual central limit theorems. The kernel $k_{1}$ (derived from $k$ and the distribution of $(\mathbf{X}^{(2)},\ldots\mathbf{X}^{(r)})$) is of order 1 and (amongst others) plays a central role in determining limiting variances in central limit theorems for U-statistics. The following central limit theorem is due to Denker and Keller (1983, Theorem 1c). We restrict attention to the case of bounded kernels, which we exclusively deal with in this paper. Boundedness could be replaced by a moment condition at the cost of strengthening the mixing condition.
@@
The asymptotic Hoeffding decomposition (A.3) explains the formula for the asymptotic variance $\sigma_{U}^{2}$. It has the form of the asymptotic variance of the process $\{k_{1}(\mathbf{X}_{i})\}_{i\in\mathbb{Z}}$, i.e. the variance of the sample mean (scaled by $\sqrt{n}$) of this process, which is a long-run variance matrix in the time series case. Note that if the kernel $k_{1}$ is degenerate, i.e. $\mathrm{Var}(k_{1}(X_{i}))=0$, the limiting distribution becomes degenerate, so a stronger scaling factor than $\sqrt{n}$ is needed to obtain a non-degenerate limiting distribution. But this is not relevant for our paper as the kernels showing up here are only degenerate in pathological cases, see Hoeffding (1948); Dengler (2010) for a discussion in the case of $\widehat{\tau}$.
There is only a small literature on variance estimation for U-statistics: For the iid case, see Shirahata and Sakamoto (1992), Wang and Lindsay (2014) and Iwashita and Klar (2024, Section 4); for the time series case the only paper that we are aware of is Dehling et al. (2017), which focuses on U-statistics of order two.
We finally provide a multivariate central limit theorem, which is the basis for many results in this paper and which follows from the univariate version in Proposition A.2 by the Cramér-Wold device and a combinatorial argument. We consider a vector of U-statistics and allow them to depend on different underlying processes. For a multivariate central limit theorem for U-statistics in the iid case see Hoeffding (1948, Theorem 7.1).
@@
Let $\{(X_{1,i},\ldots,X_{K,i})\}_{i\in\mathbb{Z}}$ be a $K$-dimensional process, where all components satisfy Assumption 1 or Assumption 2. Let $\mathbf{U}=(U^{(1)}_{\mathbf{X}_{I^{(1)}}}(k^{(1)}),\ldots,U^{(D)}_{\mathbf{X}_{I^{(D)}}}(k^{(D)}))^{\prime}$ be a $D$-dimensional vector of U-statistics with bounded symmetric kernels $\mathbf{k}=(k^{(1)},\ldots,k^{(D)})^{\prime}$ of orders $(r^{(1)},\ldots,r^{(D)})^{\prime}$ and underlying stochastic processes $\{(\mathbf{X}_{I^{(1)},i},\ldots,\mathbf{X}_{I^{(D)},i})\}_{i\in\mathbb{Z}}$, where $I^{(d)}\subset\{1,\ldots,K\}$ for $d=1,\ldots,D$ are subsets of the original index set. Furthermore, let $\theta^{(d)}$ and $k_{1}^{(d)}$, $d=1,\ldots,D$, denote the quantities defined in (A.1) and (A.4), respectively, for the $d$-th U-statistic, and define $\mbox{$\theta$}:=(\theta^{(1)},\ldots,\theta^{(D)})^{\prime}$. Then it holds that
@@
$\sigma_{vw}=\mathrm{E}\big[k^{(v)}_{1}(\mathbf{X}_{I^{(v)},i})\,k^{(w)}_{1}(\mathbf{X}_{I^{(w)},i})\big],$
@@
$\sigma_{vw}=\sum_{h=-\infty}^{\infty}\mathrm{E}\big[k^{(v)}_{1}(\mathbf{X}_{I^{(v)},i})\,k^{(w)}_{1}(\mathbf{X}_{I^{(w)},i+h})\big].$
@@
$U^{(d)}_{\mathbf{X}_{I^{(d)}}}(k^{(d)})=\binom{n}{r^{(d)}}^{-1}\sum_{1\leq i_{1}<\ldots<i_{r^{(d)}}\leq n}k^{(d)}\left(\mathbf{X}_{I^{(d)},i_{1}},\ldots,\mathbf{X}_{I^{(d)},i_{r^{(d)}}}\right)$